[Hallc_running] [New Logentry] Hall C RC update: 10/15/2021

gaskelld at jlab.org gaskelld at jlab.org
Fri Oct 15 08:35:02 EDT 2021

Logentry Text:
We will have a brief Hall C RC meeting at 10 am this morning:

Bluejeans link: https://bluejeans.com/861439920

We had pretty smooth data taking for Day shift yesterday and Owl shift today - there was no beam during part of swing shift due to the CHL trip yesterday at around 3 pm.  On day shift, there was a spike in the HMS Q2 insulating vacuum. During the accelerator down yesterday, an access was made to investigate - the local gauge in the hall indicated good vacuum, so no action was taken.

The plan for the weekend is continued production on LH2 at 42 uA with intermittent 10 cm dummy running.  A pass change is planned for Monday (Hall A to 2 pass, Hall B to 1 pass).

There will be no RC meetings over the weekend.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3925261
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211015/b3068b13/attachment.html>

More information about the Hallc_running mailing list